# African Data Ethics: A Discursive Framework for Black Decolonial Data Science

### v1

@@ 4.7. Prioritize Education & Youth
Youth Empowerment & Protection. Africa is a young continent with a large population of educated and digitally native youth (Nwankwo and Sonna, 2019; Goffi, 2023). Prioritizing the youth of Africa is a two-pronged principle: 1) protect young people from harm and 2) empower youth to lead data science agendas. The youngest generation has a tech-savviness that can be transferable to data science (African Union, 2024; Birhane, 2020; Abebe et al., 2021). If youth are expected to be the first adopters of African DDS on a large scale then these systems should be designed to protect youth so they cannot be taken advantage of. DDS should enrich the development of African youth and empower them to innovate, imagine, and contribute to bettering the communities they are a part of. Their comfort with technology may lead them to uncritically adopt a “move fast and break things” approach (Abebe et al., 2021; Ruttkamp-Bloem, 2023). To address these concerns, data science work should be intergenerational.
5. Comparative Analysis with Particularist Data Ethics Frameworks
@@
We compare our framework to six particularist frameworks to gauge the breadth of our work alongside: intersectional feminists (Klein and D’Ignazio, 2024), Western technology powers (Floridi et al., 2018), Indigenous communities of Turtle Island (Carroll et al., 2021), policymakers from global superpowers (Jobin et al., 2019), Central Asian data scientists (Younas and Zeng, 2024), and Muslim data scientists (Raquib et al., 2022). The first author reviewed each framework and used the minor principles of the African data ethics framework as a codebook (see Appendix C for comparative coding results).
@@
Each framework covered, at most, 65% of our proposed 23 principles. Delving into the diverging and converging principles provides insight into where African data ethics fits in the global data ethics discourse. The six frameworks did not discuss centering remote and rural communities or prioritization of education & youth. Our framework may have highlighted these communities because they represent a significant portion of Africa and have unique needs not fully met by the status quo (Sanny et al., 2023; Barrett et al., 2017).444https://www.theeastafrican.co.ke/tea/news/east-africa/police-army-block-ugandan-opposition-headquarters-4698696 In line with African philosophical conceptions of particularism, every community has distinct shared experiences that inform their values and normative understandings of the world. Engaging with data ethics from different cultural standpoints exposes data scientists to approaches or potential harms they would have never considered (Sadi Adamu, 2021).
@@
Common good was the only principle covered by all the frameworks. This finding also falls in line with a popular understanding of philosophy: universalism. An example of universal theory is natural law. Natural law is the legal manifestation of universalism and asserts that there are certain rights afforded to every human being(Gordon, 2022). The variety of charters, committees, and trials led by the international legal ecosystem are guided by a commitment to upholding natural law universalism. The whole RDS community is broadly guided by a universalist commitment to social good (Floridi et al., 2018). However, upon closer review, references to the common good in non-African frameworks often remain abstract, lacking the depth and specificity seen in the work of African data ethicists. African appeals to universalism are grounded in the tradition of speaking truth to power, emphasizing the need to address systemic inequities perpetuated by RDS practitioners. Drawing from critiques of international human rights law, African scholars highlight the importance of pairing narrative restoration—the recognition and reclamation of African humanity—with material restoration, including tangible reparative actions (Biko, 2004; Oyowe, 2014; Gordon, 2022). African critiques of universalism informs why African data ethicists engage with the concept of social good in ways that are both urgent and deeply pragmatic.
6. African RDS Case Studies: An Exercise of Framework Application
The primary aim of our framework is to translate its normative values and guidelines into an effective moral decision-making procedure for RDS. Moral decision-making procedures enable individuals to assess a range of choices and select actions that align with intended outcomes while upholding core values (Smith, 2022). Our framework is oriented toward advancing effective and RDS practices in Africa. The major principles describe the overarching values that should be upheld throughout the whole data science lifecycle (community benefit, challenging exploitation, and prioritizing African agency). The minor principles outline the normative practices to yield intended results and maintain the outlined values. For example, to Uphold the Universal Good in African RDS, our framework recommends proactively instituting humane work practices (Universal Dignity), evaluating the collective impact of technology (Common Good), and encouraging balance in RDS design and practices (Harmony). With our framework, we can evaluate how decisions made in African data science contexts align with the principles derived from our thematic analysis. The following section discusses two case studies from data science collectives in Africa to demonstrate how our framework can be used to evaluate responsible decision-making. To view more case study discussions, see Appendix B.
6.1. The Promise of African-led Data Science
Masakhane is an African natural language processing (NLP) collective that builds language datasets and models in Indigenous African languages. By all reports across documents, Masakhane practices all the principles proposed in this framework and especially upholds the major principle Data Self-Determination (Eke et al., 2023c; Chan et al., 2021; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023). They uphold these principles with a commitment to centering African values in their founding principles, working with existing public datasets to not infringe on data privacy, and building language datasets to preserve indigenous African languages for the future to come (Adelani et al., 2022). Masakhane also has a very welcoming and communal organizational structure to include any interested party in weekly meetings and communications. Once a member wants to contribute to a Masakhane project, they must undergo in-house training to maintain quality of the their dataset and model. They also explicitly prohibit “parachute research from the Global North” to ensure the time and resources of their collective provide direct benefit to their communities.555https://www.masakhane.io/ Finally, every project plan includes a discussion of data privacy considerations to guide their work. While there are many other parts to Masakhane’s work, the practices described in their public documents demonstrate an African data science community closely aligned with the perspectives represented in our framework. Masakhane is part of the growing grassroots efforts to imagine what African-led responsible data science can achieve.
6.2. Limits of Inclusive Representation in African Facial Recognition Technology
African people, and Black people in general, are severely underrepresented in facial recognition datasets which has led to a performance bias against Black users. African technology companies are committed to addressing this bias so their primarily African users can rely on their products. For example, a woman-led facial recognition start-up called BACE curated a diverse dataset from the local community so their facial recognition system could better detect Black subjects (Eke et al., 2023a). Users upload photos of their IDs and short videos on their phones to confirm their identity.666https://www.bacegroup.com/ For accessibility, they developed a secure mobile API. The biometric technology was first developed to aid financial fraud investigation efforts in Ghana. The current government efforts were hampered by the lack of citizens with formal identification documents. Their use cases have expanded to include identity verification for the public and private sectors. While the facial recognition technology developed by BACE achieves privacy and inclusive performance, the principles described in our framework call for more direct and sustained community involvement. Inclusion in datasets is not the only part of the data science lifecycle communities should be involved in. Additionally, BACE was created to meet the needs of financial institutions but did the team gauge if the local community had privacy or surveillance concerns during the development process (Birhane, 2020)? The work of BACE is monumental, however, our framework raises important considerations about the sustained involvement of community members.
@@ 7. Limitations
Our reflexive thematic analysis should not be conflated with a systematic literature review. While we did cast a wide net to build our corpus (including gray literature) there are impactful African data ethics documents not included in our framework. Therefore, the resulting framework should not be considered comprehensive but an introduction to notable ideas. Future work will aim to incorporate more African data ethics documents.
@@ 8. Conclusion
Through a thematic analysis of 47 documents, we derived an African data ethics framework that encompasses seven major principles: 1) decolonize & challenge internal power asymmetry, 2) center all communities, 3) uphold universal good, 4) communalism in practice, 5) data self-determination, 6) invest in data institutions & infrastructures and 7) prioritize education & youth. Our framework scratches the surface of African data ethics discourse, and the surface is rich with historically grounded, communitarian, and pragmatic insights for RDS.
A comparative analysis of our framework with six other data ethics frameworks highlights African perspectives as progressive and needed voices in global data ethics discourse. For truly pluralistic and responsible data science, we urge the data ethics community to readily seek the perspectives of Africans and other practitioners of the Global Majority. Such inclusion will not only enrich the theoretical foundation of data ethics but can inform more equitable and culturally responsive approaches to data governance, algorithmic fairness, and technological development.
Finally, a discussion of two case studies demonstrates the utility of our framework for evaluating RDS practices. The growing interest in developing and adopting AI tools for use in African contexts highlights the potential for our framework to be integrated into real-world data science workflows. In view of this, future work will engage African practitioners to evaluate the merits, gaps, and usability of our framework.
@@ Acknowledgements.
Thank you to Joseph Asike, George Obiado, Innocent Obi Jr., and Andrew Shaw for additional advice and support. The work presented was funded in part by the NSF (award # 2141506), the Sloan Foundation, and Google.References
@@
- Sanny et al. (2023) Josephine Appiah-Nyamekye Sanny, Shannon van Wyk-Khosa, and Joseph Asunka. 2023. Africa’s youth: More educated, less employed, still unheard in policy and development. Afrobarometer 734 (Nov. 2023).
@@
- Smith (2022) Holly M Smith. 2022. Making Morality Work By Holly M. Smith. Analysis 81, 4 (02 2022), 729–731. doi:10.1093/analys/anaa071 arXiv:https://academic.oup.com/analysis/article-pdf/81/4/729/42495323/anaa071.pdf
@@
- Stahl et al. (2023) Bernd Carsten Stahl, Tonii Leach, Oluyinka Oyeniji, and George Ogoh. 2023. AI policy as a response to AI ethics? Addressing ethical issues in the development of AI policies in North Africa. In Responsible AI in Africa: Challenges and opportunities. Springer International Publishing Cham, 141–167.

### v3

@@ 4.5. Center Communities on the Margins
Protect & Empower Youth. Africa is a young continent with a large population of educated and digitally native youth (Nwankwo and Sonna, 2019; Goffi, 2023). Prioritizing the youth of Africa is a two-pronged principle: 1) protect young people from harm and 2) empower youth to lead data science agendas. The youngest generation has a tech-savviness transferable to data science skills (African Union, 2024; Birhane, 2020; Abebe et al., 2021). If youth are expected to be the first adopters of African data technology then these systems should be designed to protect youth so they cannot be taken advantage of. Their comfort with technology may lead them to uncritically adopt a “move fast and break things” approach (Abebe et al., 2021; Ruttkamp-Bloem, 2023). To address these concerns, data science work should be intergenerational. African data technology should enrich the development of African youth and empower them to innovate, imagine, and contribute to bettering the communities they are a part of.
4.6. Uphold Common Good
Ethical development and deployment of DDT requires a commitment to upholding fundamental human dignity and ensuring these technologies benefit all.
Adopt Technology with a Measured Mindset. Various scholars in our corpus want the development of African data science ecosystems to be balanced, measured, inclusive, and community-minded (Kohnert, 2022; Eke et al., 2023a; Ade-Ibijola and Okonkwo, 2023). Without this approach, the adoption of AI and other data technologies can lead to more unrest and inequality across Africa. To many, the potential of DDT is profound and would change the trajectory of African development (African Union, 2024; Mabe, 2007). Data are viewed as the driving resource for the Fourth Industrial Revolution (Carman and Rosman, 2023; Gwagwa, 2019). There are African data scientists and governments who insist joining the AI boom will provide Africa the quality of life benefits afforded to the major players of past industrial revolutions (Okolo et al., 2023; Kohnert, 2022; Coetzee, 2004a).
However, there is skepticism towards wholeheartedly diving into large-scale data science adoption (Uzomah et al., 2023; Olojede, 2023). There is a need to quell the AI hype as the solution for all African problems and consider who will actually be served: Africans or the external powers propelling the AI boom (Wareham, 2021; Birhane, 2020; Sinha and Zulfa, 2023). Through the paradigm of measured development, technical development should move at the pace of social development (Kiemde and Kora, 2022; Nyerere, 1962). Paulin Hountondji’s critique of science in Africa applies well to data science development. Development should not be driven by “scientific extroversion” or catching up with global superpowers (Hountondji, 2004; Goffi, 2023). Rather, the development of data science should be an investment in the progress of African people based on African intellect, priorities, and visions of the future (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023; Biko, 2004).
Preserve the Dignity of Data Contributors. Every human and community deserves humane treatment, and African scholars in our corpus do not want data technology to ever violate human dignity (Olojede, 2023; Mhlambi, 2020). The African Charter on Human and Peoples’ Rights and the Universal Declaration on Human Rights set the precedent for the just treatment of humans (African Union, 2024). Regardless of these laws, African philosophies necessitate respect for human dignity because humans should be inherently valued for their existence and connection to others (Segun, 2021; Metz, 2021; Dignum, 2023). Every human must be treated with respect, care, and concern for their well-being (Wiredu, 2004a; Dieng, 2023; Coetzee, 2004b; Gyekye, 2004; Wiredu, 2004b). In applying the principle of universal dignity to RDS practices, every person involved in the data science lifecycle should be respected. Individuals should not be used as a means to execute data work (Metz, 2021; Ramose, 2004b). Rather, all efforts should be taken to ensure their well-being and dignity are preserved when asked to contribute to DDT (Gwagwa, 2019; Abebe et al., 2021). This same respect also extends to communities. Collective agreements need to be honored, and collective work or resources should not be used in a manner that threatens the well-being of the community (Moahi, 2007). While this principle is self-evident, there are many cases in which the rights of Africans were violated for large-scale DDT (African Union, 2024; Segun, 2021; Kohnert, 2022; Moahi, 2007).
Strive for Common Good with Systemic Change. African scholars view responsible data science as contributing to the safety, health, and goodness of all (Olojede, 2023). In various African philosophies, a person is defined by their commitment to acting for the benefit of those around them (African Union, 2024; Nwankwo and Sonna, 2019; Mhlambi and Tiribelli, 2023; Coetzee, 2004b; Ruttkamp-Bloem, 2023; Gyekye, 2004; Abdul, 2023; Wiredu, 2004b). DDT have to be made with the explicit goal of improving society and dismantling systemic harms (Sinha and Zulfa, 2023; Okolo et al., 2023; Eke et al., 2023a). In Africa, improving the efficacy of agriculture practices, healthcare access, responsiveness of public services, and the security of financial services are over-arching priorities (Carman and Rosman, 2023; Kohnert, 2022). Achieving common good involves incorporating collective values early in the process (Langat et al., 2020; Dignum, 2023), guiding development with regulatory toolkits (Olojede, 2023), not focusing on individualistic profit maximization (Gwagwa, 2019; Segun, 2021; Mabe, 2007; Nyerere, 1962; Mhlambi, 2020; Dieng, 2023), and encouraging the open sharing of data (Gwagwa, 2019; Abebe et al., 2021; Day and Rennie, 2023). The ultimate goal of RDS in Africa is upholding common good (Carman and Rosman, 2023; Metz, 2021; Olojede, 2023; Mabe, 2007).
Maintain Harmony with Stakeholders. Scholars in our corpus urge for DDT to further the mutual well-being of all stakeholders. In addition, data standards and frameworks will be most effective when they harmonize with each other (Gwagwa, 2019; Kiemde and Kora, 2022; Wareham, 2021; Mabe, 2007). In many African philosophies, harmony is not a state but a dynamic and reciprocal process of calibrating one’s actions in response to changes in the environment. In Ubuntu ethics, dogmatism is rejected because it impedes individuals from acting in harmony with the changing world (Ramose, 2004a). In Akan philosophy, morality is defined as acting in line with collective human interests (Wiredu, 2004b). Upholding harmony in data science can be understood through two dimensions: impact and practice.
Data should be harnessed to bring people closer to their environment so they can act in the best interests of not only themselves but also those around them. In terms of practice, data ethics frameworks are most effective when all the elements of data science work are accounted for (Kiemde and Kora, 2022; Gwagwa, 2019). Also, acknowledging the unique ethical needs at each stage of the data science lifecycle can inform an adaptable practice of RDS. As Gwagwa, et al. assert, the harmonious practice of RDS in Africa requires country-level data ethics frameworks to be in alignment with frameworks developed at the continental level (Gwagwa et al., 2022).
5. Orienting African Data Ethics
Now, we orient the African data ethics framework we proposed in the previous section by placing it in conversation with larger data ethics and African philosophy discourses. First, we summarize moral imperatives raised in our framework by connecting political analyses of African “development”. Next, we code data ethics framework papers from different data science communities to explore how our framework relates to the larger data ethics discourse. Finally, we highlight how our framework raises a new perspective on what it means to operationalize data ethics.
5.1. The Role of “Development” in African Data Ethics
Rethinking power relations is the root of all of the values presented in our framework. Each principle offers two-fold analyses of how power relations in African data science hamper responsible practices. On one hand, neocolonial forces bring harm to Africans in the data science lifecycle and assert control that robs Africans of agency over their DDT. On the other hand, there are internal power asymmetries in existing sociopolitical dynamics that push fellow Africans to the periphery and can lead to further marginalization. By confronting these two dimensions of power asymmetries, African data ethics shows how RDS is not only an approach to technical work but an avenue for progressive power redistribution.
The focus on self-organizing power seamlessly connects to classic African scholarship about development in Africa. Scholars such as Walter Rodney in How Europe Underdeveloped Africa presents the European extraction of African resources, exploitation of African people and suppression of African state-building as a means to keep Africa in a state of limited development (Rodney, 1982; Ndlovu-Gatsheni, 2015). However, other scholars acknowledge the impact of colonialism and imperialism but place the responsibility for Africa’s lack of development on African leaders (Gumede, 2015). The elitism, corruption, and power-hungry tendencies of African leaders harm everyday citizens and the sustainable progress of their nations (ILMI, 2014). Rather than choosing a side in this debate, our framework discourse demonstrates both problems must be overcome concurrently for progress to be made.
Many scholars go a step further to even problematize the concept of “development” because the metrics for being “developing” or “developed” are set by global superpowers (Murrey and Daley, 2023). Therefore, to imagine responsible African data science beyond AI readiness metrics, it’s important for African scholars to define what success looks like to them. Our analysis of the African data ethics corpus surfaces a set of clear goals for future African data science grounded in moral clarity and ideological traditions. These value-laden objectives guide and continue to guide the critiques, initiatives, and policies Africa develops for RDS. For examples of African data science initiatives aligned with the principles of our framework, see Appendix B.
5.2. Comparing Across Particularist Data Ethics Frameworks
@@
We compare our framework to seven particularist frameworks to gauge the breadth of our work alongside: intersectional feminists (Klein and D’Ignazio, 2024), Western technology powers (Floridi et al., 2018), Indigenous communities of Turtle Island (Carroll et al., 2021), policymakers from global superpowers (Jobin et al., 2019), Central Asian data scientists (Younas and Zeng, 2024), Muslim data scientists (Raquib et al., 2022) and Black American data scientists (Monroe-White, 2021). We call these works particularist frameworks because the authors of each paper aim to present the data science values of distinct global communities and apply the ethical theories of distinct philosophical traditions. The first author reviewed each framework and used the minor principles in our African data ethics framework as a codebook (see Appendix C for comparative coding results).
Each framework covered, at most, 68% of our proposed 19 principles. Delving into the diverging and converging principles provides insight into how African data ethics deepens global data ethics discourse through particularism.
@@
None of the frameworks discussed Centering Remote and Rural communities. Our framework highlights RDS for rural communities because they represent a significant portion of Africa and have unique needs not fully met by the status quo (Barrett et al., 2017). Communities have distinct shared experiences that inform their values and normative understandings of the world. Engaging with data ethics from different cultural standpoints exposes data scientists to approaches or potential harms they would have never considered (Sadi Adamu, 2021).
@@
Strive for Common Good with Systemic Change and Ground Design in Community-Engaged Consensus were the only principles discussed by all the frameworks. This finding also falls in line with a popular understanding of philosophy: universalism. The whole RDS community is broadly guided by a universalist commitment to social good (Floridi et al., 2018). However, upon closer review, references to the common good in non-African frameworks often remain abstract, lacking the depth and specificity seen in the work of African data ethicists. African appeals to universalism are grounded in the tradition of speaking truth to power. Drawing from critiques of international human rights law, African scholars highlight the importance of pairing narrative restoration—the recognition and reclamation of African humanity and agency—with material restoration, including tangible reparative actions (Biko, 2004; Oyowe, 2014; Gordon, 2022). Our framework corpus notes numerous examples of data science practices in Africa violating the knowledge, ownership, and safety of African communities. Social good is often taken for granted as an obvious universal goal for RDS. However, African data ethics urges the RDS community to confront the harm they have caused before it can truly achieve social good.
5.3. Reframing the Operationalization of Data Ethics
The primary critique of data ethics frameworks is the difficulty of operationalizing ideals into effectual protocols or technical practices (Munn, 2023). We offer a new perspective on the utility of data ethics frameworks: constructing frameworks is an essential operation in itself for RDS.
Articulating values and moral convictions are the first steps to contextualizing responsible data science in the epistemic traditions implicitly embedded in current practices, the socio-political status quo, and the future worlds to build for. Surfacing these contexts is important because they afford data scientists the ability to make technical decisions, develop projects, and outline long-term goals with a clear vision. The African scholars in our corpus recognize that if they don’t set the moral vision for data science in their communities, African data science will implement the divergent and often times destructive vision of global superpowers. For example, Western longtermism imagines a world in which DDT dominates humanity (Gebru and Torres, 2024), while Afrofuturism imagines a future in which DDT joins with humanity to build a more just world (Abdul, 2023). Articulating optimistic, collectivist, and resolutely African moral visions such as Afrofuturism is crucial for Africa as an emerging site of data science. Thus, as responsible data scientists, we must take responsibility for envisioning and implementing futures for ourselves and our communities. Otherwise, we remain at risk of perpetuating the injustices we seek to mitigate with RDS.
@@ 6. Limitations
Our reflexive thematic analysis should not be conflated with a systematic literature review. While we cast a wide net to build our corpus (including gray literature), there are impactful African data ethics documents not included in our framework. Therefore, the resulting framework should not be considered comprehensive but an introduction to notable ideas. Future work will aim to incorporate more African data ethics documents and engage African practitioners to evaluate the merits, gaps, and usability of our framework.
@@ 7. Conclusion
Through a thematic analysis of 50 documents, we derived an African data ethics framework that encompasses six major principles: 1) Challenge Power Asymmetries, 2) Assert Data Self-Determination, 3) Invest in Local Data Institutions & Infrastructures, 4) Utilize Communalist Practices, 5) Center Communities on the Margins, and 6) Uphold Common Good. Our framework scratches the surface of African data ethics discourse, and the surface is rich with historically grounded, communitarian, and pragmatic insights for RDS.
A comparative analysis of our framework with seven other data ethics frameworks highlights African perspectives as progressive and needed voices in global data ethics discourse. For truly pluralistic and responsible data science, we urge the RDS community to readily seek the moral perspectives of Africans, other practitioners of the Global Majority, and the particular data science communities to which they belong. Such reflexivity will not only enrich the theoretical foundation of data ethics but can inform more equitable and culturally responsive approaches to data governance, algorithmic fairness, and technological development.
@@ Acknowledgements.
Thank you to Joseph Asike, George Obaido, Innocent Obi Jr., and Andrew Shaw for additional advice and support. The work presented was funded in part by the NSF (award # 2141506), the Sloan Foundation, and Google.References
@@
- Benjamin (2019) Ruha Benjamin. 2019. Race After Technology: Abolitionist Tools for the New Jim Code. Polity, USA.
@@
- Brevini et al. (2024) Benedetta Brevini, Irene Fubara-Manuel, Clément Le Ludec, Jakob Linaa Jensen, Andrea Jimenez, and Jo Bates. 2024. Critiques of Data Colonialism (1 ed.). Bristol University Press, 120–137. http://www.jstor.org/stable/jj.15617032.11
- BROWNE (2015) SIMONE BROWNE. 2015. Dark Matters: On the Surveillance of Blackness. Duke University Press. http://www.jstor.org/stable/j.ctv11cw89p
@@
- Coleman and Batiste (2024) Grisha Coleman and Stephanie L Batiste. 2024. Blk as Tek — echo::system and Black Performance Technologies. In Proceedings of the 9th International Conference on Movement and Computing (Utrecht, Netherlands) (MOCO ’24). Association for Computing Machinery, New York, NY, USA, Article 1, 9 pages. doi:10.1145/3658852.3658854
- Cunningham et al. (2023) Jay Cunningham, Gabrielle Benabdallah, Daniela Rosner, and Alex Taylor. 2023. On the Grounds of Solutionism: Ontologies of Blackness and HCI. ACM Trans. Comput.-Hum. Interact. 30, 2, Article 20 (April 2023), 17 pages. doi:10.1145/3557890
@@
- Gebru and Torres (2024) Timnit Gebru and Émile P. Torres. 2024. The TESCREAL bundle: Eugenics and the promise of utopia through artificial general intelligence. First Monday 29, 4 (Apr. 2024). doi:10.5210/fm.v29i4.13636
@@
- Gumede (2015) Vusi Gumede. 2015. Exploring Thought Leadership, Thought Liberation and Critical Consciousness for Africa’s Development. Africa Development / Afrique et Développement 40, 4 (2015), 91–111. http://www.jstor.org/stable/90000045
@@
- ILMI (2014) AHMED ALI ILMI. 2014. CHAPTER EIGHT: Rethinking Development: An Indigenous African Communal Approach. Counterpoints 443 (2014), 142–156. http://www.jstor.org/stable/42982053
@@
- Monroe-White (2021) Thema Monroe-White. 2021. Emancipatory Data Science: A Liberatory Framework for Mitigating Data Harms and Fostering Social Transformation. In Proceedings of the 2021 Computers and People Research Conference (Virtual Event, Germany) (SIGMIS-CPR ’21). Association for Computing Machinery, New York, NY, USA, 23–30. doi:10.1145/3458026.3462161
@@
- Munn (2023) Luke Munn. 2023. The uselessness of AI ethics. AI and Ethics 3, 3 (2023), 869–877.
- Murrey and Daley (2023) Amber Murrey and Patricia Daley. 2023. Beyond Tokenism: Pluriversals and Decolonizing Solidarity for Thriving and Dignified Futures (1 ed.). Pluto Press, 203–227. http://www.jstor.org/stable/jj.5938918.12
@@
- Ndlovu-Gatsheni (2015) Sabelo J. Ndlovu-Gatsheni. 2015. Genealogies of Coloniality and Implications for Africa’s Development. Africa Development / Afrique et Développement 40, 3 (2015), 13–40. http://www.jstor.org/stable/afrdevafrdev.40.3.13
- Noble (2018) Safiya Noble. 2018. Algorithms of Oppression: How Search Engines Reinforce Racism. NYU Press, USA.
@@
- Reece et al. (2025) Shaniah Reece, Covenant Adenuga, Tierra Ablorh, and Yolanda A. Rankin. 2025. A Powerful Sisterhood: Black Women Scholars in Feminist HCI. Interactions 32, 1 (Jan. 2025), 44–47. doi:10.1145/3705739
@@
- Rodney (1982) Walter Rodney. 1982. How Europe Underdeveloped Africa. Howard University Press.